OpenStudy (anonymous):

In a geometric sequence, if a2=162 and a5=-6, what is the value of r??

OpenStudy (anonymous):

a5/a2 = r^3 hence r^3= -6/162 or r^3= -1/27 or r^3=(-1/3)^3 hence r=-1/3

OpenStudy (anonymous):

why do you do that?

OpenStudy (anonymous):

"a5/a2 = r^3"

OpenStudy (anonymous):

omg! i was subtracting instead of dividing T_T!! thanks so much!

OpenStudy (anonymous):

see a5= (a1)*(r^4) and a2=(a1)*(r) then a5/a2 =((a1)*(r^4))/((a1)*(r))=r^4/r=r^3

OpenStudy (anonymous):

i get it! :) ty